a) if the price elasticity of demand for labor is 0.1 and the wage increased from $10 to $15 an hour, what is the predicted decrease in the level of employment in percentage terms?


Instructions : Enter your response as a whole number.



B) if the price elasticity of demand for labor is 0.2 and the wage increased from $10 to $15 an hour, what is the predicted decrease in the level of employment in percentage terms?


Instructions: around your response to one decimal place.
Answer #1

Elasticity of demand for labor = % Change in Demand for Labor/ % Change in wage

Elasticity of demand for labor = 0.1

% Change in wage = (15-10)/10*100 = 50%

Elasticity of demand for labor = % Change in Demand for Labor/ % Change in wage

0.1 = % Change in Demand for Labor/50

% Change in Demand for Labor = 50*0.1 = 5 Percent

Elasticity of demand for labor = % Change in Demand for Labor/ % Change in wage

Elasticity of demand for labor = 0.2

% Change in wage = (15-10)/10*100 = 50%

Elasticity of demand for labor = % Change in Demand for Labor/ % Change in wage

0.2 = % Change in Demand for Labor/50

% Change in Demand for Labor = 50*0.2 = 10 Percent